# Fine-Tuning or Retrieval? Comparing Knowledge Injection in LLMs

### 10 Dec 2023 v1

@@ Abstract
Large language models (LLMs) encapsulate a vast amount of factual information within their pre-trained weights, as evidenced by their ability to answer diverse questions across different domains. However, this knowledge is inherently limited, relying heavily on the characteristics of the training data. Consequently, using external datasets to incorporate new information or refine the capabilities of LLMs on previously seen information poses a significant challenge. In this study, we compare two common approaches: fine-tuning and retrieval-augmented generation (RAG). We evaluate both approaches on a variety of knowledge-intensive tasks across different topics. Our findings reveal that while fine-tuning offers some improvement, RAG consistently outperforms it, both for existing knowledge encountered during training and entirely new knowledge. Moreover, we find that LLMs struggle to learn new factual information through fine-tuning, and that exposing them to numerous variations of the same fact during training could alleviate this problem.
@@ 4 Knowledge Base Creation
Orca2 7B 0.578 0.652 0.593 0.674
Astronomy (0-shot) Mistral 7B 0.625 0.684 0.651 0.697
Llama2 7B 0.421 0.461 0.487 0.520
Orca2 7B 0.651 0.763 0.651 0.75
Astronomy (5-shot) Mistral 7B 0.658 0.704 0.651 0.691
Llama2 7B 0.388 0.474 0.447 0.520
Orca2 7B 0.664 0.770 0.664 0.743
@@
Llama2 7B 0.340 0.410 0.360 0.390
Orca2 7B 0.450 0.470 0.370 0.380
(a) Results for the MMLU datasets described in Section 4.1 in terms of log-likelihood accuracy (Equation 4).
@@
(b) Current events results. Models that were fine-tuned on the original dataset are labeled as FT-reg, while those trained on the dataset with multiple paraphrases are labeled as FT-par.
@@ 4.1 Task Selection and Rationale
MMLU Benchmark  To properly evaluate the capabilities of LLMs on knowledge-intensive tasks, we selected four distinct tasks from the Massively Multilingual Language Understanding Evaluation (MMLU) benchmark (Hendrycks et al., 2021) in the topics of anatomy, astronomy, college biology, and college chemistry. The chosen tasks were selected based on their emphasis on factual knowledge and the minimal reliance on reasoning. As a heuristic, we opted for tasks where the questions are short and involve no context. This approach aims to enable us to test LLM proficiency in comprehending and manipulating information in isolation from its reasoning processes.
@@ 5 Experiments and Results
Figure 2: Box plot comparing all knowledge-injection methods over all experiments in  Table 0(a).
@@
MMLU Results  For each task and model, we compared four approaches: using just the base model, RAG, FT, and finally combining FT and RAG by using the fine-tuned model as the generator. Furthermore, we tested the MMLU tasks using both 0-shot and 5-shot scenarios. The full results are shown in  Table 0(a) and summarized in  Figure 2.
In all cases, RAG performed significantly better compared to the base models. Furthermore, using RAG with the base model as the generator was consistently better than only fine-tuning. In some cases, using the fine-tuned model instead of the base model as the generator in the RAG pipeline can improve results even further. However, this is not consistent and demonstrates the instability of fine-tuning. Additionally, we found that the 5-shot approach boosts the results by a small margin in most cases, although not consistently.
Current Events Results  The evaluation on the current events task is shown in Table 0(b). RAG proves particularly effective due to the one-to-one correspondence between the questions and the auxiliary dataset (see Section 4.3). Fine-tuning is not competitive with RAG. However, fine-tuning with multiple paraphrases still provides a significant improvement over the baseline. We note that combining RAG with fine-tuning shows inferior performance compared to RAG alone.
It is worth noting that although the questions are based on information the models were not exposed to during training, the results of the base models surpass $\frac{1}{L}=0.25$. We deduce that this is due to the correlation between past and present events, along with the high reasoning capabilities of Orca2-7B and Mistral-7B.
@@ Appendix A RAG Ablation Study
Llama2 7B 0.511 0.563 0.533 0.526 0.519
Orca2 7B 0.622 0.652 0.600 0.607 0.600
Astronomy (0-shot) Mistral 7B 0.651 0.664 0.658 0.664 0.684
Llama2 7B 0.461 0.454 0.447 0.447 0.461
Orca2 7B 0.704 0.730 0.737 0.763 0.724
Astronomy (5-shot) Mistral 7B 0.697 0.645 0.664 0.664 0.704
Llama2 7B 0.467 0.474 0.461 0.441 0.461
Orca2 7B 0.730 0.704 0.750 0.750 0.770
@@
Llama2 7B 0.493 0.521 0.493 0.479 0.472
Orca2 7B 0.618 0.639 0.632 0.653 0.660
Chemistry (0-shot) Mistral 7B 0.490 0.480 0.450 0.420 0.430
Llama2 7B 0.360 0.370 0.430 0.410 0.370
Orca2 7B 0.400 0.400 0.440 0.450 0.380
Chemistry (5-shot) Mistral 7B 0.480 0.440 0.380 0.430 0.410
Llama2 7B 0.360 0.380 0.390 0.330 0.320
Orca2 7B 0.430 0.400 0.460 0.440 0.430

### 30 Jan 2024 v3

@@ Abstract
Large language models (LLMs) encapsulate a vast amount of factual information within their pre-trained weights, as evidenced by their ability to answer diverse questions across different domains. However, this knowledge is inherently limited, relying heavily on the characteristics of the training data. Consequently, using external datasets to incorporate new information or refine the capabilities of LLMs on previously seen information poses a significant challenge. In this study, we compare two common approaches: unsupervised fine-tuning and retrieval-augmented generation (RAG). We evaluate both approaches on a variety of knowledge-intensive tasks across different topics. Our findings reveal that while unsupervised fine-tuning offers some improvement, RAG consistently outperforms it, both for existing knowledge encountered during training and entirely new knowledge. Moreover, we find that LLMs struggle to learn new factual information through unsupervised fine-tuning, and that exposing them to numerous variations of the same fact during training could alleviate this problem.
Keywords: LLMs, NLP, Fine-Tuning vs. RAG, Knowledge and Factuality.
@@ 4 Knowledge Base Creation
Table 1: Results for the MMLU datasets described in Section 4.1 in terms of log-likelihood accuracy (Equation 4).
@@
Orca2 7B 0.570 0.659 0.593 0.674
Astronomy (0-shot) Mistral 7B 0.625 0.678 0.651 0.697
Llama2 7B 0.401 0.467 0.487 0.520
Orca2 7B 0.645 0.750 0.651 0.750
Astronomy (5-shot) Mistral 7B 0.658 0.724 0.651 0.697
Llama2 7B 0.401 0.474 0.447 0.520
Orca2 7B 0.664 0.763 0.664 0.743
@@
Llama2 7B 0.370 0.380 0.360 0.390
Orca2 7B 0.430 0.470 0.370 0.380
Prehistory (0-shot) Mistral 7B 0.713 0.750 0.719 0.731
Llama2 7B 0.448 0.481 0.457 0.478
Orca2 7B 0.642 0.679 0.673 0.673
Prehistory (5-shot) Mistral 7B 0.722 0.762 0.725 0.762
Llama2 7B 0.515 0.531 0.503 0.537
Orca2 7B 0.664 0.698 0.667 0.694
Table 2: Current events results. Models that were fine-tuned on the original dataset are labeled as FT-reg, while those trained on the dataset with multiple paraphrases are labeled as FT-par.
@@ 4.1 Task Selection and Rationale
MMLU Benchmark  To properly evaluate the capabilities of LLMs on knowledge-intensive tasks, we selected four distinct tasks from the Massively Multilingual Language Understanding Evaluation (MMLU) benchmark (Hendrycks et al., 2021) in the topics of anatomy, astronomy, college biology, college chemistry and prehistory. The chosen tasks were selected based on their emphasis on factual knowledge and the minimal reliance on reasoning. As a heuristic, we opted for tasks where the questions are short and involve no context. In practice we selected four STEM subjects as well as one humanities subject, to ensure the evaluation is not limited to certain fields. Note that prehistory involves questions spanning all non-modern history. This approach aims to enable us to test LLM proficiency in comprehending and manipulating information in isolation from its reasoning processes.
@@ 5 Experiments and Results
Figure 2: The relative accuracy gain (as explained in Equation 5) for each knowledge-injection method, averaged (columnwise) across all experiments in Table 1.
@@
MMLU Results  For each task and model, we compared four approaches: using just the base model, RAG, FT, and finally combining FT and RAG by using the fine-tuned model as the generator. Furthermore, we tested the MMLU tasks using both 0-shot and 5-shot scenarios. The full results are shown in  Table 1. An aggregation of the relative accuracy gain, i.e.,
$(\mathcal{L}_{\mathcal{M^{\prime}},\mathcal{Q}}-\mathcal{L}_{\mathcal{M},\mathcal{Q}})/{\mathcal{L}_{\mathcal{M},\mathcal{Q}}},$ (5)
where $\mathcal{M}$ is the base model and $\mathcal{M^{\prime}}$ is the knowledge-injected model, is shown in  Figure 2.
In all cases, RAG performed significantly better compared to the base models. Furthermore, using RAG with the base model as the generator was consistently better than only fine-tuning. In some cases, using the fine-tuned model instead of the base model as the generator in the RAG pipeline improved results even further. However, this is not consistent and thus demonstrates the inherent instability of fine-tuning. Additionally, we found that the 5-shot approach boosts the results by a small margin in most cases, with a similar trend being observed in all of the different approaches.
Current Events Results  The evaluation on the current events task is shown in Table 2. RAG proves particularly effective due to the one-to-one correspondence between the questions and the auxiliary dataset (see Section 4.3). Fine-tuning is not competitive with RAG. However, fine-tuning with multiple paraphrases still provides a significant improvement over the baseline. We note that combining RAG with fine-tuning shows inferior performance compared to RAG alone.
It is worth noting that although the questions are based on information the models were not exposed to during training, the results of the base models surpass $\frac{1}{L}=0.25$. This can partially be explained by the models using reasoning and/or pre-existing knowledge when answering questions that are not independent of the past information. Some examples of this can be found in Appendix C.
@@ Appendix A RAG Ablation Study
Llama2 7B 0.496 0.563 0.541 0.526 0.526
Orca2 7B 0.630 0.659 0.600 0.600 0.600
Astronomy (0-shot) Mistral 7B 0.651 0.678 0.678 0.664 0.664
Llama2 7B 0.447 0.434 0.447 0.434 0.467
Orca2 7B 0.711 0.730 0.730 0.750 0.730
Astronomy (5-shot) Mistral 7B 0.704 0.684 0.658 0.684 0.724
Llama2 7B 0.461 0.447 0.474 0.428 0.454
Orca2 7B 0.730 0.737 0.750 0.743 0.763
@@
Llama2 7B 0.500 0.521 0.507 0.465 0.472
Orca2 7B 0.625 0.639 0.625 0.660 0.660
Chemistry (0-shot) Mistral 7B 0.450 0.470 0.470 0.500 0.470
Llama2 7B 0.320 0.320 0.300 0.380 0.360
Orca2 7B 0.370 0.420 0.400 0.410 0.440
Chemistry (5-shot) Mistral 7B 0.540 0.490 0.500 0.510 0.470
Llama2 7B 0.280 0.320 0.340 0.340 0.380
Orca2 7B 0.390 0.430 0.400 0.430 0.470
Prehistory (0-shot) Mistral 7B 0.728 0.725 0.750 0.735 0.728
Llama2 7B 0.481 0.460 0.457 0.457 0.429
Orca2 7B 0.648 0.645 0.660 0.670 0.679
Prehistory (5-shot) Mistral 7B 0.710 0.750 0.759 0.756 0.762
Llama2 7B 0.512 0.485 0.525 0.519 0.531
Orca2 7B 0.660 0.688 0.685 0.698 0.688
@@ Appendix B Paraphrase Examples
Appendix C Current Events Existing Knowledge Examples
To give a better understanding of how a model might be able to answer questions about new information, with better than random success, we present three possible scenarios as examples. These scenarios show how models with stronger reasoning skills can infer the correct answer even for unseen information.
The first scenario involves questions about previously unseen information, where basic reasoning abilities allow a model to make an educated guess.
Question: What was a key issue that led to the 2023 United Auto Workers strike?
Answers:
  1. 1.
    Dissatisfaction with the quality of cafeteria food.
  2. 2.
    Disagreements over employee dress codes.
  3. 3.
    Discontent with stagnant wages and tiered employment systems.
  4. 4.
    Debates over the color scheme of the factories.
In this case it is easy to guess that the third option is the most likely, even without knowledge of this specific strike.
A second scenario involves questions where prior knowledge about a topic may aid a model in answering.
Question: What environmental concern was raised by some scientists as a result of the 2023 Hawaii wildfires?
Answers:
  1. 1.
    Rising temperatures.
  2. 2.
    Melting ice caps.
  3. 3.
    Charred soils running off into the shoreline.
  4. 4.
    Increased air pollution.
In this case, knowing the geography of Hawaii, as well as immediate effects of wildfires, enables a model to give the first two options a lower likelihood. This process of elimination increases the probability of choosing one of the remaining options (the third option is the correct answer).
A third scenario arises due to the automatic question generation process, some questions strongly rely on pre-existing knowledge.
Question: What event in 2021 was compared to the September 2023 New York floods?
Answers:
  1. 1.
    Hurricane Katrina.
  2. 2.
    Hurricane Ida.
  3. 3.
    Hurricane Sandy.
  4. 4.
    Hurricane Harvey.
Since only one of these events occurred in 2021 (Hurricane Ida), and all the models tested have been exposed to events from 2021 during pre-training, this question can potentially be answered without using additional current information.
Finally, to demonstrate why it is reasonable to assume that models cannot generally answer questions about new information, with better than random success, look at the following example:
Question: How did Matthew Belk, a National Weather Service meteorologist, describe the September 2023 northeastern U.S. floods?
Answers:
  1. 1.
    50-year event.
  2. 2.
    100-year event.
  3. 3.
    200-year event.
  4. 4.
    500-year event.
Even with some knowledge about floods and their statistical properties, it would be very difficult to guess that this specific meteorologist would call the flood a ‘200-year event’. This is especially true if the model was not exposed to information about the details of the flood.